The cost of equipment purchased by Charleston, Inc., on June 1, 2020, is $89,000. It is estimated that the machine will havePlease provide me all explaination in detail with formulas?
Answer #1

A. Straight line method.

>> Depreciation per annum = ( $ 89,000 - $ 5,000 ) / 7 = $ 12,000

>> Depreciation for 2020 = $ 12,000 * ( 7 /12 ) = $ 7,000.

>> Depreciation for 2021 = $ 12,000.

________________________________________________________________________________________

B. Units of output.

>> Depreciation per unit = ( $ 89,000 - $ 5,000 ) / 525,000 = $ 0.16.

>> Depreciation for 2020 = 55,000 * $ 0.16 = $ 8,800.

>> Depreciation for 2021 = 48,000 * $ 0.16 = $ 7,680.

__________________________________________________________________________________________

C. Working ours.

>> Depreciation per hours = ( $ 89,000 - $ 5,000 ) / 42,000 = $ 2

>> Depreciation for 2020 = 6,000 * $ 2 = $ 12,000.

>> Depreciation for 2021 = 5,500 * $ 2 = $ 11,000.

______________________________________________________________________________________________

D. Sum of digits method

>> Sum of years = 7 ( 7 +1 ) / 2 = 28.

>> Depreciation for 2020 = ( $ 89,000 - $ 5,000 ) * ( 7 / 28 ) * ( 7 / 12 ) = $ 12,250.

>> Depreciation for 2021 = ( $ 89,000 - $ 5,000 ) * ( 6.41667 / 28 ) = $ 19,250.

___________________________________________________________________________________________

E . Double declining balance

>> Depreciation rate = 200 / 7 = 28.57 %.

>> Depreciation for 2020 = $ 89,000 * 28.57 % * ( 7 /12 ) = $ 14,833.

>> Depreciation for 2021 = ( $ 89,000 - $ 14,833 )* 28.57 % = $ 21,190

_____________________________________________________________________________________________